OpenStudy (mtalhahassan2):

A taxi cab drives 2.0 km [W], then 3.0 km [N], then 4.0 km [W], and finally 5.0 km [N]. The entire trip takes 0.30 h. What is the taxi's average velocity? Question 4 options: A) 33 km/h [53° N of W] B) 47 km/h [53° W of N] C) 10 km/h [53° N of W] D) 33 km/h [53° W of N] E) 47 km/h [53° N of W]

OpenStudy (priyar):

average velocity = total displacement / total time taken the " total time taken" is directly given to you.. So find the displacement ...for finding that its better to draw the diagram for this path... so that we can find the initial and final points..

OpenStudy (michele_laino):

total displacement is: 

OpenStudy (michele_laino):

sorry, I have made a typo: 

OpenStudy (michele_laino):

namely: \[d = \sqrt {{8^2} + {6^2}} = 10\;Km\]

OpenStudy (michele_laino):

and the direction, is: \[\theta = \arccos \left( {\frac{3}{5}} \right) = ...?\]